# Beyond the Surface: Enhancing LLM-as-a-Judge Alignment with Human via Internal Representations

### v1

@@ Abstract
The growing scale of evaluation tasks has led to the widespread adoption of automated evaluation using large language models, a paradigm known as "LLM-as-a-judge." However, improving its alignment with human preferences without complex prompts or fine-tuning remains challenging. In this work, motivated by preliminary findings that middle-to-upper layers encode semantically and task-relevant representations that are often more aligned with human judgments than the final layer, we propose lager, a lightweight and efficient framework for enhancing LLM-as-a-Judge alignment with human scoring, via internal representations. lager produces fine-grained judgment scores by aggregating cross-layer score-token logits and computing the expected score from a softmax-based distribution, with the LLM backbone kept frozen. lager fully leverages the complementary information across different layers, overcoming the limitations of relying solely on the final layer. We evaluate our method on the standard alignment benchmarks Flask, HelpSteer, and BIGGen using Spearman correlation, and find that lager achieves improvements of up to 7.5% over the best baseline across these benchmarks. Without reasoning steps, lager matches or outperforms reasoning-based methods. Experiments on downstream applications, such as data selection and emotional understanding, further show the effectiveness of our method.
@@ 1 Introduction
In this paper, we propose an effective LLM-as-a-judge framework called lager, to enhance LLM-as-a-Judge alignment with human via internal representations. Inspired by prior studies (Wang et al., 2020; Yang et al., 2022; Roh et al., 2024) showing that middle-to-upper layers encode richer semantic and task-specific information, we propose to aggregate the logits of score tokens from different layers. By applying a weighted combination followed by a softmax, we obtain a probability distribution over candidate scores, from which the final score is computed as the expected value. lager leverages the semantic diversity across different layers, allowing the final score to integrate both low-level lexical cues and high-level reasoning signals, thereby providing a more comprehensive reflection of the model’s understanding of the scoring task. Unlike the finetuning-based method, the parameters of the general LLM backbone are always kept frozen. lager outperforms baselines on alignment benchmarks such as Flask, HelpSteer, and BIGGen, with improvements of up to 7.5% in Spearman correlation. Although reasoning-based evaluation methods provide more transparency by exposing intermediate reasoning steps, their reasoning is often shallow and fails to improve reliability, leading to inferior performance. In contrast, without explicit reasoning, lager achieves comparable or superior results to these reasoning-based baselines. We further select instruction data with lager method and achieve better performance on the AlpacaEval-2.0 benchmark than various baselines. Experiments on more downstream applications show the effectiveness of the lager method.  [^1] Our code will be publicly available at https://github.com/sustech-nlp/LAGER.
@@ 2 Preliminary
When incorporating a general LLM as the judge for a model response, the prompt template includes the evaluation task description $x_{d}$, user instruction $x_{i}$, the response to be evaluated $x_{r}$, and the scoring criteria and output format requirement $x_{c}$ (Liu et al., 2023, 2024; Hu et al., 2024a). We denote the full input as $X=\{x_{d},x_{i},x_{r},x_{c}\}$, and the set of candidate scores as $\mathbb{S}\subset\mathcal{V}$, where $\mathcal{V}$ is the model’s vocabulary. Let $\mathbf{h}^{(L)}_{n}$ denote the final-layer hidden state at the position where the score token is to be generated (e.g., after ‘Score:’). The model produces output logits:
@@ 3.2 Score Estimation Based on LLM Hidden Representations
$s^{*}=\mathbb{E}_{s\sim P(s)}[s]=\sum_{s\in\mathbb{S}}s\times P(s)$ (12)
@@ 3.3 Lightweight Training of Layer Weights
We propose two types of layer weights $\bm{w}$. One is to apply average aggregation $w_{l}=1/(L+1)$ (denoted as lager (w.o. tuning) in Table 1), the other is to tune the lightweight $L+1$ parameters with a small-scale validation set (denoted as lager (w. tuning)). With a frozen backbone and minimal learnable parameters, we refer to the dataset as a validation set to distinguish it from finetuning-based LLM evaluators. To enhance the model’s performance while aligning its predictions more closely with the distribution of human scores, we adopt a combination of cross-entropy(CE) loss and mean absolute error (MAE) loss, balanced by a weighting hyperparameter $\alpha$.
@@
where $\mathcal{B}$ is the batch of samples in a training step, and $s_{truth}^{i}$ is the human-annotated score for each sample. Please refer to Appendix E.5 for specific training settings and details. It is important to note that the weights are tuned only once for each backbone and subsequently applied across all benchmarks. As the LLM remains frozen, no additional judge model is needed for scenarios like self-improvement. In contrast, finetuning-based methods require domain-specific judge models and suffer from limited generalization due to scarce labeled data.
@@ 3.4 Discussion: Applicability under Restricted Access Settings
While our proposed framework relies on accessing the intermediate layers of the LLM, it can naturally adapt to more restricted scenarios. For instance, when only final-layer logits are available (e.g., API-based models), we can extract the logits corresponding to score tokens, apply softmax, and compute the expected score. The method proposed in prior work (Liu et al., 2023) generates a continuous score estimate by avoiding argmax on discrete tokens. We refer to this setup as expectation score (E-Score). The evaluation uses the vanilla score, based on the most likely token, when only the final token prediction is observable. Both settings can be viewed as simplified special cases within our framework. We treat both scores as comparable baselines in our experiments.
@@ 4.1 Experiments Setup
Benchmarks
Due to potential preference leakage (Li et al., 2025) between models, we evaluate our method on human-annotated datasets for reliable results. We chose three point-wise benchmarks to ensure a comprehensive evaluation: Flask (Ye et al., 2024a), HelpSteer (Wang et al., 2023), and BiGGen Bench [^3] https://huggingface.co/datasets/prometheus-eval/BiGGen-Bench-Results  (Kim et al., 2024b). In these benchmarks, human-annotated scores range from 1 to 5. Please refer to Appendix E.2 for the details.
Models
We experiment with multiple backbone models of varying sizes and from different families, including Mistral-7B-Instruct-v0.3 (Jiang et al., 2023a), InternLM3-8B-Instruct (Cai et al., 2024), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), Qwen-2.5-14B-Instruct (Qwen et al., 2025), Mistral-Small-24B-Instruct and LLAMA-3.3-70B-Instruct.
Baselines
We compare lager with three baselines: GPTScore, vanilla score (VScore) and expectation score (E-Score). At the same time, we compare the commonly used API-based model GPT-4o-mini. We also experiment with methods that explicitly train the LLM backbones, namely TIGERScore-7B and Prometheus2-7B. While lager does not modify the backbone models, these methods are included solely for reference and are not directly compared. Please refer to Appendix E.3 for specific details about the baselines.
Evaluation Metrics
Following previous works (Gu et al., 2024; Bai et al., 2023; Liu et al., 2025), we use the commonly adopted Pearson and Spearman correlations to measure the consistency between LLM-predicted scores and human annotations. See Appendix E.4 for further details.
Evaluation details
We adopt greedy decoding for its stability and consistency in evaluation (Lin & Chen, 2023). To comprehensively assess the effectiveness of our method, we conduct evaluations under two different conditions: 1) Direct: Directly providing a score during evaluation without reasoning. 2) Reasoning: Previous work (Liu et al., 2023; Wang et al., 2024a; Hu et al., 2024a) shows that reasoning improves judgment quality, with reasoning-first setups performing best(Chen et al., 2024; Murugadoss et al., 2024; Gu et al., 2024). Thus, we adopt a reasoning-first approach in this study. Please refer to Appendix G for the prompt templates. Unless otherwise specified, we highlight the best results in each table using bold, and denote the second-best with underlining.
Table 1: Spearman Correlation Results (Part 1 of 2): lager vs. Baselines on Flask, HelpSteer, and BiGGen Bench. See Table 5 for the remaining results. These results are statistically significant, with p-values less than 1e-5. See Table 6 for the full Pearson correlation results.
@@ 4.2 Main Results
Tables 1 and 5 report Spearman results of lager compared to baselines across three benchmarks. Section 5.3 further analyzes the performance variation of lager and baseline methods across different model scales, under the setting where models from the Qwen2.5 family are used as frozen backbones, and compares their results on the same set of benchmarks.
The Effectiveness and Scalability of LAGER
lager with tuning consistently outperforms all non-training baselines across diverse LLM backbones on all three benchmarks. This highlights the importance of leveraging internal logits and output distributions for accurate scoring. Although Mistral-Small-24B-Instruct and Qwen2.5-14B-Instruct underperform GPT-4o-mini on Vscore, they achieve comparable results when using lager. GPTScore performs poorly due to its reliance on generation probabilities, which overlook the semantic and logical quality. Due to its focus on error detection, the fine-tuned TIGERscore-7B shows relatively limited performance. Compared to the fine-tuned Prometheus2-7B, lager also enables InternLM3-8B-Instruct and LLaMA3.1-8B-Instruct(results shown in Table 5), two models of comparable size, to surpass it even though they initially have weaker performance on Vscore.
The Robustness of LAGER with Tuning
The average E-Score performance of models like Qwen-2.5-14B-Instruct, LLAMA-3.3-70B-Instruct, and Mistral-7B-Instruct-v0.3 is sometimes lower than VScore. Specifically, the average E-score performance of LLAMA-3.3-70B-instruct (results shown in Table 5) across all benchmarks is 2.4 points lower than Vscore. In contrast, lager with tuning achieves almost the best performance across all settings and benchmarks, consistently outperforms Vscore. lager can closely match human score distribution and incorporate more fine-grained information. More details are discussed in Section 5.
Reasoning Is Not Always Better
The richness of information aggregation increases gradually from Vscore, E-Score to lager (with tuning). Figure 6 shows that reasoning leads to an improvement in Vscore performance, a divergence in E-Score performance (with half of the models improving and the other half declining), and consistently weaker results for lager (with tuning) under the reasoning setting compared to direct. We speculate that this phenomenon may be due to reasoning potentially causing the model to become overconfident (Yang et al., 2024; Wagner et al., 2024), which in turn leads to inaccurate evaluation scores.
4.3 Ablation Study
@@
To better understand each design choice of lager, we conduct a comprehensive ablation study on the direct setup using Mistral-7B-Instruct-v0.3 and InternLM3-8B-Instruct models. Table 2 shows the results of the Spearman correlation. The expectation (Exp.) score and the maximum (Max.) score are two alternative designs. Given the score distribution of a certain layer, we use Eqn.12 to calculate the expected score or use the score with the highest probability.
Logits agg. and prob. agg. are two alternative designs, aggregating information from all hidden layers at the logits level or output distribution level. Removing logits agg. and prob. agg. indicates that we do not use internal states. It is observed that the full configuration (Exp. score + Logits agg. + Tuning) achieves the best performance across all settings, with a maximum Spearman correlation of 0.545; Fine-tuning brings up to +0.10 improvement; expectation scoring outperforms maximum scoring (up to +0.17); and logits aggregation surpasses probability aggregation (up to +0.07). Multi-layer integration is generally effective, especially for Mistral; however, in untuned settings, InternalLM occasionally performs better without aggregation, indicating model-specific differences.
@@ 5.2 LAGER Yields More Human-Aligned and Fine-Grained Score Distributions
We also quantitatively compare lager, E-Score and VScore by evaluating the Kullback-Leibler (KL) divergence and the mean squared error (MSE). The KL divergence of a candidate score $S$ is computed as $\mathrm{D}_{\mathrm{KL}}(\tilde{p}_{hs}||\tilde{p}_{s})$ where $\tilde{p}_{hs}$ and $\tilde{p}_{s}$ represent the human score distribution and the candidate score distribution, respectively. The results, presented in Table 3, demonstrate that lager more accurately approximates the distribution of human-annotated scores.

### v3

@@ Abstract
The growing scale of evaluation tasks has led to the widespread adoption of automated evaluation using LLMs, a paradigm known as "LLM-as-a-judge". However, improving its alignment with human preferences without complex prompts or fine-tuning remains challenging. Previous studies mainly optimize based on shallow outputs, overlooking rich cross-layer representations. In this work, motivated by preliminary findings that middle-to-upper layers encode semantically and task-relevant representations that are often more aligned with human judgments than the final layer, we propose lager, a post-hoc, plug-and-play framework for improving the alignment of LLM-as-a-Judge point-wise evaluations with human scores, by leveraging internal representations. lager produces fine-grained judgment scores by aggregating cross-layer score-token logits and computing the expected score from a softmax-based distribution, while keeping the LLM backbone frozen and ensuring no impact on the inference process. lager fully leverages the complementary information across different layers, overcoming the limitations of relying solely on the final layer. We evaluate our method on the standard alignment benchmarks Flask, HelpSteer, and BIGGen using Spearman correlation, and find that lager achieves improvements of up to 7.5% over the best baseline across these benchmarks. Without reasoning steps, lager matches or outperforms reasoning-based methods. Experiments on downstream applications, such as data selection and emotional understanding, further show the generalization of lager.
@@ 1 Introduction
In this paper, we propose an effective LLM-as-a-judge framework called lager, to enhance LLM-as-a-Judge alignment with human via internal representations. Inspired by prior studies (Wang et al., 2020; Yang et al., 2022; Roh et al., 2024) showing that middle-to-upper layers encode richer semantic and task-specific information, we propose to aggregate the logits of score tokens from different layers. By applying a weighted combination followed by a softmax, we obtain a probability distribution over candidate scores, from which the final score is computed as the expected value. lager leverages the semantic diversity across different layers, allowing the final score to integrate both low-level lexical cues and high-level reasoning signals, thereby providing a more comprehensive reflection of the model’s understanding of the scoring task. lager is plug-and-play, leaving model parameters and next-token predictions unchanged. lager outperforms baselines on alignment benchmarks such as Flask, HelpSteer, and BIGGen, with improvements of up to 7.5% in Spearman correlation. Although reasoning-based evaluation methods provide more transparency by exposing intermediate reasoning steps, their reasoning is often shallow and fails to improve reliability, leading to inferior performance. In contrast, without explicit reasoning, lager achieves comparable or superior results to these reasoning-based baselines. We further select instruction data with lager method and achieve better performance on the AlpacaEval-2.0 benchmark than various baselines. Experiments on more downstream applications show the effectiveness of the lager method.  [^1] Our code is publicly available at https://github.com/sustech-nlp/LAGER.
@@ 2 Preliminary
When incorporating a general LLM as the judge for a model response, the prompt template includes the evaluation task description $x_{d}$, user instruction $x_{i}$, the response to be evaluated $x_{r}$, and the scoring criteria and output format requirement $x_{c}$ (Liu et al., 2023, 2024; Hu et al., 2024a). We denote the full input as $X=\{x_{d},x_{i},x_{r},x_{c}\}$, and the set of candidate scores as $\mathbb{S}\subset\mathcal{V}$, where $\mathcal{V}$ is the model’s vocabulary. Let $\mathbf{h}^{(L)}_{n}$ denote the final-layer hidden state at the position where the score token is to be generated (e.g., after ’Score:’). The model produces output logits:
@@ 3.2 Score Estimation Based on LLM Hidden Representations
$s^{*}=\mathbb{E}_{s\sim P(s)}[s]=\sum_{s\in\mathbb{S}}s\times P(s).$ (12)
@@ 3.3 Lightweight Training of Layer Weights
We propose two types of layer weights $\bm{w}$. One is to apply average aggregation $w_{l}=1/(L+1)$ (denoted as lager (w.o. tuning) in Table 1), the other is to tune the lightweight $L+1$ parameters ($L$ is the number of transformer layers) with a small-scale validation set (denoted as lager (w. tuning)). With a frozen backbone and minimal learnable parameters, we refer to the dataset as a validation set to distinguish it from finetuning-based LLM evaluators. To enhance the model’s performance while aligning its predictions more closely with the distribution of human scores, we adopt a combination of cross-entropy(CE) loss and mean absolute error (MAE) loss, balanced by a weighting hyperparameter $\alpha$.
@@
where $\mathcal{B}$ is the batch of samples in a training step, and $s_{truth}^{i}$ is the human-annotated score for each sample. Please refer to Appendix E.5 for specific training settings and details. It is important to note that the weights are tuned only once for each backbone and subsequently applied across all benchmarks and downstream tasks. As the LLM remains frozen, no additional judge model is needed for scenarios like self-improvement. Our method is implemented without altering the model’s next-token prediction process or logits, thereby offering plug-and-play capability. In contrast, finetuning-based methods require domain-specific judge models and suffer from limited generalization due to scarce labeled data.
@@ 3.4 Discussion: Applicability under Restricted Access Settings
While lager relies on accessing the intermediate layers of the LLM, it can naturally adapt to more restricted scenarios. For instance, when only final-layer logits are available (e.g., API-based models), we can extract the logits corresponding to score tokens, apply softmax, and compute the expected score (this method is proposed in G-Eval (Liu et al., 2023)). We refer to this setup as expectation score (E-Score). The evaluation uses the vanilla score, based on the most likely token, when only the final token prediction is observable. Both settings can be viewed as simplified special cases within our framework. We treat both scores as comparable baselines in our experiments. Nonetheless, powerful open-source models still remain the mainstream for large-scale evaluation due to their lower cost and performance comparable to closed-source models (Lambert et al., 2024; Gu et al., 2024; Malik et al., 2025), making lager broadly applicable and easily extensible to future models.
@@ 4.1 Experiments Setup
Benchmarks. Due to potential preference leakage (Li et al., 2025) between models, we evaluate our method on human-annotated datasets for reliable results. We chose three diverse point-wise benchmarks to ensure a comprehensive evaluation: Flask (Ye et al., 2024a), HelpSteer (Wang et al., 2023), and BiGGen Bench [^3] https://huggingface.co/datasets/prometheus-eval/BiGGen-Bench-Results  (Kim et al., 2024b). In these benchmarks, human-annotated scores range from 1 to 5. Please refer to Appendix E.2 for the details.
Models. We experiment with multiple backbone models of varying sizes and from different families, including Mistral-7B-Instruct-v0.3 (Jiang et al., 2023a), InternLM3-8B-Instruct (Cai et al., 2024), LLaMA3.1-8B-Instruct (Grattafiori et al., 2024), Qwen-2.5-14B-Instruct (Qwen et al., 2025), Mistral-Small-24B-Instruct and LLAMA-3.3-70B-Instruct.
Baselines. We compare lager with three baselines: GPTScore, vanilla score (VScore) and expectation score (E-Score). At the same time, we compare the commonly used API-based model GPT-4o-mini. We also experiment with methods that explicitly train the LLM backbones, namely TIGERScore-7B and Prometheus2-7B. While lager does not modify the backbone models, these methods are included solely for reference and are not directly compared. Please refer to Appendix E.3 for specific details about the baselines.
Evaluation Metrics. Following previous works (Gu et al., 2024; Bai et al., 2023; Liu et al., 2025), we use the commonly adopted Pearson and Spearman correlations to measure the consistency between LLM-predicted scores and human annotations. See Appendix E.4 for further details.
Table 1: Spearman Correlation Results: lager vs. Baselines on Flask, HelpSteer, and BiGGen Bench. These results are statistically significant, with p-values less than 1e-5. See Table 5 for the full Pearson correlation results.
@@
Mistral-7B-Instruct-v0.3
GPTScore 0.258 - 0.209 - 0.183 - 0.217
Vscore 0.266 0.269 0.267 0.364 0.138 0.280 0.264
E-Score 0.239 0.279 0.296 0.380 0.185 0.283 0.277
lager (w.o tuning) 0.338 0.295 0.401 0.377 0.353 0.329 0.349
lager (w. tuning) 0.347 0.298 0.403 0.376 0.357 0.333 0.352
LLaMA3.1-8B-Instruct
GPTScore 0.061 - -0.022 - -0.162 - -0.041
Vscore 0.334 0.429 0.374 0.518 0.273 0.390 0.386
E-Score 0.386 0.446 0.464 0.525 0.352 0.403 0.429
lager (w.o tuning) 0.472 0.456 0.520 0.524 0.475 0.443 0.482
lager (w. tuning) 0.477 0.460 0.515 0.524 0.482 0.444 0.484
@@
LLAMA-3.3-70B-Instruct
GPTScore 0.042 - 0.014 - -0.193 - -0.046
Vscore 0.518 0.567 0.435 0.494 0.559 0.539 0.519
E-Score 0.464 0.540 0.444 0.488 0.530 0.506 0.495
lager (w.o tuning) 0.610 0.598 0.506 0.520 0.597 0.585 0.569
lager (w. tuning) 0.611 0.598 0.504 0.519 0.602 0.584 0.570
Evaluation Details. We adopt greedy decoding for its stability and consistency in evaluation (Lin & Chen, 2023). To comprehensively assess the effectiveness of our method, we conduct evaluations under two different conditions: 1) Direct: Directly providing a score during evaluation without reasoning. 2) Reasoning: Previous work (Liu et al., 2023; Wang et al., 2024a; Hu et al., 2024a) shows that reasoning improves judgment quality, with reasoning-first setups performing best(Chen et al., 2024; Murugadoss et al., 2024; Gu et al., 2024). Thus, we adopt a reasoning-first approach in this study. The prompt template used in our work follows the standard configuration widely adopted in prior studies (Kim et al., 2024b; Liu et al., 2023; Xu et al., 2023), please refer to Appendix G for the prompt templates. Unless otherwise specified, we highlight the best results in each table using bold, and denote the second-best with underlining.
@@ 4.2 Main Results
Tables 1 report Spearman results of lager compared to baselines across three benchmarks. Section 5.3 further analyzes the performance variation of lager and baseline methods across different model scales, under the setting where models from the Qwen2.5 family are used as frozen backbones, and compares their results on the same set of benchmarks. Specifically, lager (w. tuning) achieves near-best performance on these benchmarks and demonstrates strong stability, so it serves as the main approach in our work.
The Effectiveness and Scalability of LAGER.
lager with tuning consistently outperforms all non-training baselines across diverse LLM backbones on all three benchmarks. This highlights the importance of leveraging internal logits and output distributions for accurate scoring. Although Mistral-Small-24B-Instruct and Qwen2.5-14B-Instruct underperform GPT-4o-mini on Vscore, they achieve comparable results when using lager. GPTScore performs poorly due to its reliance on generation probabilities, which overlook the semantic and logical quality. Due to its focus on error detection, the fine-tuned TIGERscore-7B shows relatively limited performance. Compared to the fine-tuned Prometheus2-7B, lager also enables InternLM3-8B-Instruct and LLaMA3.1-8B-Instruct, two models of comparable size, to surpass it even though they initially have weaker performance on Vscore. Even without tuning, lager outperforms E-Score in 5 out of 6 backbone models, with an average improvement of 4.5%. In the few cases where it underperforms, the performance gap is within 3%. Moreover, it consistently outperforms V-Score.
The Robustness of LAGER.
The average E-Score performance of models like Qwen-2.5-14B-Instruct, LLAMA-3.3-70B-Instruct, and Mistral-7B-Instruct-v0.3 is sometimes lower than VScore. Specifically, the average E-score performance of LLAMA-3.3-70B-instruct across all benchmarks is 2.4 points lower than Vscore. In contrast, lager with tuning achieves almost the best performance across all settings and benchmarks, consistently outperforms Vscore. lager can closely match human score distribution and incorporate more fine-grained information. More details are discussed in Section 5. Therefore, we primarily recommend the lager with tuning approach and conduct subsequent analyses based on it. The performance gains of LAGER stem not from large-scale fine-tuning, but from a more efficient and structured utilization of the model’s internal representations.
Reasoning Is Not Always Better.
The richness of information aggregation increases gradually from Vscore, E-Score to lager (with tuning). Figure 6 shows that reasoning leads to an improvement in Vscore performance, a divergence in E-Score performance (with half of the models improving and the other half declining), and consistently weaker results for lager (with tuning) under the reasoning setting compared to direct. We speculate that this phenomenon may be due to reasoning potentially causing the model to become overconfident (Yang et al., 2024; Wagner et al., 2024), which in turn leads to inaccurate evaluation scores. This is also consistent with the findings of previous work (Ge et al., 2025; Guo et al., 2024; Li et al., 2024d): As the reasoning steps progress, the model’s focus on the original input and the text to be evaluated gradually diminishes, shifting more toward its self-generated reasoning trajectory.
@@
4.3 Ablation Study
To better understand each design choice of lager, we conduct a comprehensive ablation study on the direct setup using Mistral-7B-Instruct-v0.3 and InternLM3-8B-Instruct. Table 2 shows the results of the Spearman correlation. The expectation (Exp.) score and the maximum (Max.) score are alternative designs. Given the score distribution of a certain layer, we use Eqn.12 to calculate the expected score or use the score with the highest probability. Logits agg. and prob. agg. are two alternative designs, aggregating information from all hidden layers at the logits level or output distribution level. Removing logits agg. and prob. agg. indicates that we do not use internal states. It is observed that the full configuration (Exp. score + Logits agg. + Tuning) achieves the best performance across all settings, with a maximum Spearman correlation of 0.545; Fine-tuning brings up to +0.10 improvement; expectation scoring outperforms maximum scoring (up to +0.17); and logits aggregation surpasses probability aggregation (up to +0.07). Multi-layer integration is generally effective, especially for Mistral; however, in untuned settings, InternalLM occasionally performs better without aggregation, indicating model-specific differences.
@@ 5.2 LAGER Yields More Human-Aligned and Fine-Grained Score Distributions
We also quantitatively compare lager, E-Score, and VScore by evaluating the Kullback-Leibler (KL) divergence and the mean squared error (MSE). The KL divergence of a candidate score $S$ is computed as $\mathrm{D}_{\mathrm{KL}}(\tilde{p}_{hs}||\tilde{p}_{s})$ where $\tilde{p}_{hs}$ and $\tilde{p}_{s}$ represent the human score distribution and the candidate score distribution, respectively. The results in Table 3, demonstrate that lager more accurately approximates the distribution of human-annotated scores.